If the supply of corn by the Namibia is Q_n=a+bp and the supply by the rest of the world is Q_r=c+ep, what is the world supply? What is the world inverse supply?
1
Expert's answer
2015-06-04T00:00:45-0400
The world supply can be calculated as follows:
Q_w = Q_n + Q_r
So,
Q_w = a + bp + c + ep = a + c + p(b + e)
Then the inverse supply equation would be:
P = (Q_w – a – c)/(b + e)